OpenStudy (anonymous):

Find the solution to the system of equations represented by the matrix shown below 1 9 -1|48 -7 1 7|48 4 1 -9|-58 A.x = 8, y = 6, z = 8 B.x = 2, y = 6, z = 8 C.x = 14, y = 54, z = 56

OpenStudy (anonymous):

b.

OpenStudy (anonymous):

From Mathematica:\[\text{LinearSolve}\left[\left( \begin{array}{ccc} 1 & 9 & -1 \\ -7 & 1 & 7 \\ 4 & 1 & -9 \end{array} \right),\left( \begin{array}{c} 48 \\ 48 \\ -58 \end{array} \right)\right]\to \{\{2\},\{6\},\{8\}\} \]